OpenStudy (firekat97):

Hi, how would you integrate ∫√(3)x(2 − 4x)^(1/3). dx ?? Thanks in advance! :)

OpenStudy (irishboy123):

draw it, maybe? that could be interpreted in any number of ways...

OpenStudy (irishboy123):

or scan or link the source

OpenStudy (firekat97):

\[\int\limits \sqrt{3}x(2-4x)^{1/3}.dx\]

OpenStudy (irishboy123):

by parts u = x, v' = (2-4x)^{1/3)

OpenStudy (irishboy123):

\[\int \sqrt{3}x(2-4x)^{1/3} \ dx\] \[\sqrt{3} \int x(2-4x)^{1/3} \ dx\] \[u = x, u' = 1\] \[v' = (2-4x)^{1/3}, v = (3/4) (-1/4) (2-4x)^{4/3}\]

OpenStudy (irishboy123):

like that

OpenStudy (firekat97):

ohhh okay thanks!!